# EntGPT: Linking Generative Large Language Models with Knowledge Bases

### v1

@@
Abstract.
The ability of Large Language Models (LLMs) to generate factually correct output remains relatively unexplored due to the lack of fact-checking and knowledge grounding during training and inference. In this work, we aim to address this challenge through the Entity Disambiguation (ED) task. We first consider prompt engineering, and design a three-step hard-prompting method to probe LLMs’ ED performance without supervised fine-tuning (SFT). Overall, the prompting method improves the micro-$F_{1}$ score of the original vanilla models by a large margin, on some cases up to 36% and higher, and obtains comparable performance across 10 datasets when compared to existing methods with SFT. We further improve the knowledge grounding ability through instruction tuning (IT) with similar prompts and responses. The instruction-tuned model not only achieves higher micro-F1 score performance as compared to several baseline methods on supervised entity disambiguation tasks with an average micro-$F_{1}$ improvement of 2.1% over the existing baseline models, but also obtains higher accuracy on six Question Answering (QA) tasks in the zero-shot setting. Our methodologies apply to both open- and closed-source LLMs.
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXjournal: JACMjournalvolume: 37journalnumber: 4article: 111publicationmonth: 8ccs: Information systems Question answeringccs: Information systems Information extractionccs: Computing methodologies Information extractionccs: Computing methodologies Natural language generation
1. Introduction
@@
Figure 1. Pipeline of the proposed frameworks: Given input document with the annotated mention, both the approaches first conducts (1) entity candidate generation step to obtain relevant entities. EntGPT-Prompting (EntGPT-P) then performs (2) an augmentation step to obtain an auxiliary content of the annotated mention followed by (3) a multi-choice selection prompt to decide the corresponding entity of the annotated mention. EntGPT-Instruction tuning (EntGPT-I), on the other hand, goes through the (2) construction of the instruction dataset, and is (3) instruction-tuned on the curated data.
The introduction of Large Language Models (LLMs) has sparked intense interest and scrutiny regarding tools that leverage their capabilities. Although their text-generation and unstructured reasoning capabilities are “remarkable” (Raffel et al., 2020; Radford et al., 2018, 2019; Brown et al., 2020; Touvron et al., 2023b), their ability to produce structured output remains underdeveloped and relatively unexplored (Yu et al., 2022; Sun et al., 2023). Moreover, tools leveraging the LLMs are known to generate content that is factually incorrect and nonsensical (Ji et al., 2023). This phenomena, commonly known as hallucination, occurs, at least partly, due to a lack of factual knowledge grounding during pre-training. To address this issue, researchers have introduced various strategies such as encompassing feedback mechanisms, external information retrieval, and early refinement in language model generation (Tonmoy et al., 2024). However, these models typically do not ground the generated knowledge to well-structured knowledge bases. In the present work, we mitigate these common problems by linking (i.e., grounding) keyterms to their entries in a knowledge base and using that structured knowledge to support the text generation task.
The entity disambiguation (ED) task is a type of information extraction (IE) that seeks to link text fragments that represent some real-world entity with an entry in some structured knowledge base like an encyclopedia or a dictionary. Once linked, the existing data and its relationships within the knowledge base could be used to assist in a number of downstream processes. The integration of deep neural network models, such as Transformers (Vaswani et al., 2017), utilized in LLMs, with the symbolic reasoning capabilities of existing knowledge bases, has been dubbed as the Third Wave of AI (DARPA, 2018) by the United States Defense Advanced Research Projects Agency (DARPA). This convergence is considered essential for the advancement of problem solving and reasoning in AI systems (Brooks, 1981; Hitzler and Sarker, 2022). For such a vision to become reality, it is critical that LLMs be linked to the symbolic entities that they reference.
@@
In the present work and as illustrated in Fig. 1, we explore two distinct approaches to ground the generative language model through the ED task. The first is a multi-step prompting framework, which emulates prompt-tuning called Entity GPT-Prompting (EntGPT-P). This approach first generates a set of entity candidates for the mentions in the document. Then, it utilizes the power of the LLM to generate auxiliary content facilitating the final selection of the corresponding entity from the candidate set. Results over various tasks and datasets show that EntGPT-P achieves comparable performance to supervised models even without training/fine-tuning on human-annotated data.
@@
Secondly, we implemented an Instruction Tuning (IT) approach , named EntGPT-Instruction tuning (EntGPT-I), which is further trained on instructions (i.e., Prompt-Response pairs) and exhibits more controllable and predictable behavior (Zhang et al., 2023b). EntGPT-I obtained significantly better results (micro-$F_{1}$ score) compared to EntGPT-P across ED benchmarks. We also evaluated EntGPT-I on six closed-world question answering tasks. Experiments showed that EntGPT-I achieved state of the art (SOTA) performance (micro-$F_{1}$ score) in a zero-shot setting while boosting accuracy scores over the most popular and powerful commercial LLMs.
Our major contributions can be summarized as follows:
@@
  • We present a model instruction-tuned on generative LLMs(i.e., EntGPT-I) with (instruction - response) pairs for the ED task, while augmenting the instructions with enhanced knowledge. We achieve SOTA performance on the ED task.
@@
  • We also provide a detailed performance comparison of our proposed approaches on both open-source (e.g., Llama2) and closed-source (e.g., GPT-3.5) LLMs. We find a consistent boost in micro-$F_{1}$/accuracy performance when compared to their vanilla counterparts.
@@
2. Related Work
2.1. Hallucination in LLMs
The so-called hallucination problem has impeded the use of the intelligent systems that leverage LLMs in sensitive applications such as summarizing financial analysis reports, medical records, and customer support interactions (Tonmoy et al., 2024). To overcome the challenges raised by the generation of inaccurate and erroneous information, researchers have proposed several techniques. In a recent survey paper, Tonmoy et. al (Tonmoy et al., 2024) categorized these techniques into two different subgroups: Prompt Engineering, and Models Development. Prompt engineering explores generating more factual and grounded output with well-curated instructions/context to the model; model development requires algorithmic and/or architectural improvement.
Within the prompt engineering paradigm, researchers have explored the use of external knowledge to provide updated and more-thorough context to the model (Peng et al., 2023; Varshney et al., 2023; Lewis et al., 2020), refinements to the generated output through proper feedback to the model (Si et al., 2022; Dhuliawala et al., 2023; Lei et al., 2023; Mündler et al., 2023), and prompt tuning (Cheng et al., 2023; Jones et al., 2023). Moreover, techniques involving decoding strategies (Shi et al., 2023; Chuang et al., 2023; Li et al., 2023), use of of Knowledge Graphs (KGs) (Ji et al., 2022; Bayat et al., 2023), faithfulness-based loss functions (Yoon et al., 2022; Qiu et al., 2023), and supervised fine-tuning (Elaraby et al., 2023; Tian et al., 2023; Zhang et al., 2023a) have been introduced to tackle hallucination through the algorithmic approach. The EntGPT-P method described in the present work falls under the prompting engineering paradigm, whereas the EntGPT-I model tackles hallucination through an instruction-tuning approach.
2.2. Entity Disambiguation
@@
Existing research in ED is primarily divided into two categories: feature extraction and task refinement. Key work in feature extraction includes the interaction between mentions and context (Ganea and Hofmann, 2017; Kolitsas et al., 2018), the consistency between entities and entity types (Raiman and Raiman, 2018; Tedeschi et al., 2021; Raiman, 2022), the relationship between entities and knowledge base entries (Ayoola et al., 2022a, b), and correlation between entities (Phan et al., 2019; Xiao et al., 2023b). Traditional ED models framed this task as entity classification (Ganea and Hofmann, 2017), while more recent work approached it as machine reading comprehension (Barba et al., 2022a), dense retrieval (Wu et al., 2020), question answering (Zhang et al., 2022b), sequence-to-sequence generation (Cao et al., 2021), and prompt LLMs (Xiao et al., 2023a).
2.3. Prompt-based Methods for Information Extraction
One of the important developments in the field of prompt-based methods for information extraction was the release of GPT-3, which has been shown to outperform state-of-the-art methods for language modeling (Brown et al., 2020). The benefit of prompt-based methods is the ability to query a language model to obtain additional knowledge and understanding about a given context, thereby significantly improving performance in zero-shot and few-shot settings (Wang et al., 2023b; Sun et al., 2023).
@@ 2.3. Prompt-based Methods for Information Extraction
There are mainly three streams of work in this paradigm. The first is in-context learning where a few demonstrative examples are provided to an LLM in order to obtain some desired output. The second is prompt-tuning, where a set of learnable parameters are inserted into the model (Lester et al., 2021; Ding et al., 2022) or each layer (Li and Liang, 2021; Wang et al., 2023a) for fine-tuning. The third line of work is instruction-tuning, where (instruction, response) pairs are prepared used to fine-tune the LLM (Longpre et al., 2023).
@@
In terms of task formulation, existing work on prompt-based methods for information extraction can be primarily categorized into three paradigms: type-oriented, span-oriented, and generation. Type-oriented tasks (Ding et al., 2022) mainly locate the given class for a certain mention within the original documents. Span-oriented methods (Cui et al., 2021) enumerate the possible spans and assign corresponding class labels. Finally, generation aims to transform information extraction into a sequence generation schema (Wang et al., 2023b).
3. Methodology
@@ 3.1. Problem Definition
3.2. EntGPT-P: Entity Disambiguation with Multi-Step Prompting
@@
3.2.1. Step 1: Entity Candidate Generation
Given a mention $m$ in document $D$, the first step is to find a subset of entity candidates $\mathcal{E}_{c}$ corresponding to $m$ within the knowledge base $\mathcal{E}$. We combine two strategies to generate entity candidates. First, we use statistical information of hyperlinks in the knowledge base to surface the most likely entity candidates $\mathcal{E}_{p}$ from $\mathcal{E}$ that exhibit syntactic similarity to the mentions within the document (Ganea and Hofmann, 2017). However, this probabilistic approach, we call the Prior probability-approach or simply Prior, suffers from low recall because it only considers syntactic similarity. In low-recall cases, we additionally employ a dense retrieval model as a backup to augment the entity candidate generation. Specifically, we use the Bidirectional Encoder Representations from Transformers (BERT)-based entity linking (BLINK) model, which primarily uses cleaned Wikipedia hyperlinks to as its retrieval model (Wu et al., 2020), to generate extra entity candidates $\mathcal{E}_{r}$. Therefore, the final entity candidates set from step 1 is $\mathcal{E}_{c}=\mathcal{E}_{p}\cup\mathcal{E}_{r}$ from which we select the top-10 candidates.
3.2.2. Step 2: Augmentation by Prompting
@@
3.2.3. Step 3: Multiple-choice Selection by Prompting
@@
3.3. EntGPT-I: Entity Disambiguation with Instruction Tuning
In an initial study, described below, we found that EntGPT-P substantially reduces hallucination and helped to deliver more accurate results compared to non-augmented foundation models (see metrics in Tables 2 and  3). Nonetheless, EntGPT-P still under-performed supervised methods indicating room for improvement. To that end, we constructed an additional instruction-tuning model on top of generative LLMs through similar prompts as that of EntGPT-P. In this section, we discuss the processes involved in creating the EntGPT-I model including the construction of an instruction dataset and the instruction-tuning procedure.
3.3.1. Instruction Dataset Construction
An instruction dataset usually consists of three different elements: an instruction, an optional input, and an anticipated output (Ji et al., 2023). Usually in the form of natural language, an instruction provides the description of the task to the model, whereas any optional input can provide additional context. In the present work, we curate one of the largest manually annotated entity linking and entity disambiguation datasets, i.e., Accurate online Disambiguation of Entities in Text and Tables (AIDA) (Hoffart et al., 2011) into the required format (instruction, input, output) needed to fine-tune the an LLM model. An example of such a training instance is shown in Fig1 (at right).
3.3.2. Instruction Tuning
We fine-tune two generative models (GPT3.5 and Llama2-chat) in a fully supervised manner on the curated instruction dataset. We used the OpenAI API platform (Kublik and Saboo, 2023) to fine-tune GPT3.5, and we used Replicate API [^1] https://replicate.com/docs/get-started/python to fine-tune Llama2-chat. Fine-tuning was performed for only one epoch and we kept all other hyperparameters at their default settings. Although hyper-parameter tuning could potentially enhance performance, hyper-parameter search is outside the scope of this study. After fine-tuning, EntGPT-I is then evaluated on both the ED task and the question answering task.
4. Experiments
@@
  • RQ1: Does multi-step prompting (EntGPT-P) and instruction tuning (EntGPT-I) produce more accurate output (i.e., reduce hallucination)?
@@
4.1. Datasets
Following general ED settings introduced in prior work (Guo and Barbosa, 2018; Kolitsas et al., 2018; Ayoola et al., 2022b; Barba et al., 2022a), we evaluated our models on six ED datsets: AIDA (Milne and Witten, 2008), MSNBC (MSN) (Cucerzan, 2007), Aquaint text corpus (AQUAINT/AQU) (Milne and Witten, 2008), a multilingual training corpus ACE2004 (ACE04) (Ratinov et al., 2011), Walking Named Entity Disambiguation (WNED-WIKI) (Guo and Barbosa, 2018), and WNED-CWEB (CWEB) (Gabrilovich et al., 2013). We also consider five additional datasets from GERBIL (Usbeck et al., 2015): Keyphrase Overlap Relatedness for Entity Disambiguation (KORE50/KORE) (Hoffart et al., 2012), Open knowledge extraction challenge (OKE-2015/OKE15) (Nuzzolese et al., 2015), OKE-2016 (OKE16) (Nuzzolese et al., 2016), N3-Reuters-128 (REU) (Röder et al., 2014), and N3-RSS-500 (RSS) (Röder et al., 2014). We provide the statistics of these benchmarks, including the number of documents, and mentions in Table 1. Each of the experiments used Wikipedia as the background knowledge base. Following the example of Guo and Barbosa (2018), we manually removed spurious mentions that do not appear in the knowledge base as well as repeated documents and empty documents without any mentions.
We found that EntGPT-I (with GPT3.5 as the foundation model) outperformed other approaches on the average micro-F1 score (Tables 2 and  3), so we additionally evaluated EntGPT-I (GPT3.5) on six different commonsense reasoning (question answering) benchmarks: AI2 Reasoning Challenge-Challenge Set (ARC-C) (Clark et al., 2018), OpenbbookQA (OBQA) (Mihaylov et al., 2018), Abductive Natural Language Inference (aNLI) (Bhagavatula et al., 2019), CommonsenseQA (CSQA) (Talmor et al., 2018), Physical Interaction QA (PIQA) (Bisk et al., 2020), and Social Interaction QA (SIQA) (Sap et al., 2019).
4.2. Comparison Models

### v3

# EntGPT: Entity Linking with Generative Large Language Models

@@
Abstract
Entity Linking (EL) in natural language processing seeks to match text entities to their corresponding entries in a dictionary or knowledge base. Traditional approaches rely on contextual models, which can be complex, hard to train, and poorly transferable across domains. Generative large language models (LLMs) like GPT offer a promising alternative but often underperform with naive prompts. In this study, we introduce EntGPT, employing advanced prompt engineering to enhance EL tasks. Our three-step hard-prompting method (EntGPT-P) significantly boosts the micro-$F_{1}$ score by up to 36% over vanilla prompts, achieving competitive performance across 10 datasets without supervised fine-tuning (SFT). Additionally, our instruction tuning method (EntGPT-I) improves micro-$F_{1}$ scores by 2.1% on average in supervised EL tasks and outperforms several baseline models in six Question Answering (QA) tasks. Our methods are compatible with both open-source and proprietary LLMs. All data and code are available on GitHub at https://github.com/yifding/In_Context_EL.
keywords:
Information Extraction, Generative Large Language Models, Entity Linking
journal: Knowledge-Based Systems
[…]
[label1]organization=University of Notre Dame, city=Notre Dame, state=Indiana, country=USA
[…]
1 Introduction
@@
Figure 1: Pipeline of EntGPT frameworks: Given input document with the annotated mention, both the approaches first conducts (1) entity candidate generation step to obtain relevant entities. EntGPT-Prompting (EntGPT-P) then performs (2) an augmentation step to obtain an auxiliary content of the annotated mention followed by (3) a multi-choice selection prompt to decide the corresponding entity of the annotated mention. EntGPT-Instruction tuning (EntGPT-I), on the other hand, goes through the (2) construction of the instruction dataset, and is (3) instruction-tuned on the curated data.
The advent of Large Language Models (LLMs) has generated significant excitement and concern in the fields of natural language processing (NLP) and artificial intelligence (AI). While their capabilities in text generation and unstructured reasoning are impressive Raffel-JMLR'20-T5; Radford-OpenAI_Blog'18-GPT; Radford-OpenAI_Blog'19-GPT2; Brown-NIPS'20-GPT3; Touvron-arXiv'23-Llama2, their ability to produce structured output remains underdeveloped and relatively unexplored Zhu-AAAI'22-JAKET; Dong-arXiv'23-Head-to-Tail. Entity disambiguation (ED), a crucial task in information extraction (IE), involves linking text fragments representing real-world entities to entries in structured knowledge bases like encyclopedias or dictionaries. These links can enhance various downstream processes by leveraging existing data and relationships within the knowledge base. Integrating deep neural network models, such as Transformers Vaswani-NIPS'17-attention with the symbolic reasoning capabilities of knowledge bases has been termed the “Third Wave of AI”Darpa-Darpa'18-Next_AI. This convergence is essential for advancing problem-solving and reasoning in AI systemsBrooks-AI'81-Symbolic; Hitzler-Book'22-Neuro. For this vision to materialize, it is critical that LLMs are effectively linked to the symbolic entities they reference.
Recent research efforts Qin-arXiv'23-Chatgpt_NLP_Task; Gonzalez-arXiv'23-ChatGPT_Historical_Entities; wei2023zero; li2023prompting have utilized LLMs, specifically GPT-3.5, for IE tasks. However, these studies primarily focus on named entity recognition (NER) and relation extraction (RE) tasks having only a few dozen entity types, which is far less than the 6 million entities present in general Wikipedia entries for the entity disambiguation (ED) task. The current state-of-the-art in ED relies heavily on entity features, predicting entity labels (such as Wikipedia pages or DBpedia entries) from models trained on Wikipedia links Ganea-EMNLP'17-deep_ed, Web hyperlinks Wu-EMNLP'20-BLINK, or info-boxes Ayoola-NAACL'22-ReFinED; Bhargav-NAACL'22-Dbpedia_Entity_Type, and considering the internal structure among entities Hu-KBS'20-Graph_Entity. These methods have been successful in closed-world scenarios where data is clean and complete Hoffart-EMNLP'11-CONLL03_entity, and in open-world scenarios where data may be incomplete but well-described Logeswaran-ACL'19-zero; Wu-EMNLP'20-BLINK. Another approach to the ED task is through text generation. One of the earliest generative models for ED is GENRE DeCao-ICLR'21-GENDRE, which proposed a sequence-to-sequence framework that generates an entity-label sequence from a mention-sequence conditioned with special indicators.
In this work, we use LLMs for the ED task. Most errors in ED arise from rare entities, but auxiliary knowledge from LLMs can help address these challenging cases. Furthermore, most ED models require a ranking step to select the best entity from several candidates within a given context. Generative LLMs have shown exceptional ability in classification tasks xu2023large; fei2022lasuie; josifoski2022genie and can naturally fit into the entity ranking step with their extensive background knowledge.
@@
As illustrated in Fig. 1, we explore two distinct approaches for tackling the ED task with LLMs. The first approach is a multi-step prompting framework called EntGPT-Prompting (EntGPT-P). This method first generates a set of entity candidates for mentions in a document. Then, it uses the LLM to generate auxiliary content, aiding the final selection of the correct entity from the candidate set. Results across various tasks and datasets show that EntGPT-P achieves performance comparable to supervised models without requiring training or fine-tuning on human-annotated data.
@@
The second approach is called EntGPT-Instruction Tuning (EntGPT-I). This method is trained further on instructions (i.e., prompt-response pairs) and demonstrates more controllable and predictable behavior zhang2023instruction. EntGPT-I significantly outperformed EntGPT-P in micro-$F_{1}$ scores across ED benchmarks. Additionally, we evaluated EntGPT-I on six closed-world question-answering tasks. Experiments showed that EntGPT-I achieved state-of-the-art (SOTA) performance in the zero-shot setting, significantly improving accuracy scores over the most popular commercial LLMs.
In summary, our major contributions are as follows:
  • 1.
@@
  • 2.
    We present EntGPT-I, a model instruction-tuned on generative LLMs with instruction-response pairs for the ED task, while augmenting the instructions with enhanced knowledge. We achieve SOTA performance on the ED task.
  • 3.
@@
  • 4.
    We also provide a detailed performance comparison of both EntGPT approaches on open-source (e.g., Llama2) and closed-source (e.g., GPT-3.5) LLMs. We find a consistent boost in micro-$F_{1}$/accuracy performance when compared to their vanilla counterparts.
@@
2 Related Work
2.1 Entity Disambiguation
@@
Existing research in ED is primarily divided into two categories: feature extraction and task refinement. Key work in feature extraction includes the interaction between mentions and context Ganea-EMNLP'17-deep_ed; Kolitsas-CONLL'18-end2end, the consistency between entities and entity types Raiman-AAAI'18-deeptype; Tedeschi-EMNLP'21-NER4EL; Raiman-AAAI'22-deeptype2, the relationship between entities and knowledge base entries Ayoola-NAACL'22-kb; Ayoola-NAACL'22-ReFinED, and correlation between entities Phan-TKDE'19-Pair_link; Xiao-EMNLP'23-CoherentED. Traditional ED models framed this task as entity classification Ganea-EMNLP'17-deep_ed, while more recent work approached it as machine reading comprehension Barba-ACL'22-Extend, dense retrieval Wu-EMNLP'20-BLINK, question answering Zhang-ICLR'22-Entqa, sequence-to-sequence generation DeCao-ICLR'21-GENDRE, and prompt LLMs Xiao-EMNLP'23-Instruction.
2.2 Prompt-based Methods for Information Extraction
@@
One of the important developments in the field of prompt-based methods for information extraction was the release of GPT-3, which has been shown to outperform state-of-the-art methods for language modeling Brown-NIPS'20-GPT3. The benefit of prompt-based methods is the ability to query a language model to obtain additional knowledge and understanding about a given context, thereby significantly improving performance in zero-shot and few-shot settings Li-arXiv'23-GPT-NER; Dong-arXiv'23-Head-to-Tail. There are mainly three streams of work in this paradigm. The first is in-context learning where a few demonstrative examples are provided to an LLM in order to obtain some desired output. The second is prompt-tuning, where a set of learnable parameters are inserted into the model Lester-EMNLP'21-power; Ding-ACL'22-Prompt_Entity_Typing or each layer Li-ACL'21-prefix; Wang-EMNLP'23-aprompt for fine-tuning. The third line of work is instruction-tuning, where (instruction, response) pairs are prepared used to fine-tune the LLM Longpre-arXiv'23-flan.
@@
In terms of task formulation, existing work on prompt-based methods for information extraction can be primarily categorized into three paradigms: type-oriented, span-oriented, and generation. Type-oriented tasks Ding-ACL'22-Prompt_Entity_Typing mainly locate the given class for a certain mention within the original documents. Span-oriented methods Cui-ACL'21-template_NER_BART enumerate the possible spans and assign corresponding class labels. Finally, generation aims to transform information extraction into a sequence generation schema Li-arXiv'23-GPT-NER.
3 EntGPT
@@ 3.1 Problem Definition
3.2 EntGPT-P: Entity Disambiguation with Multi-Step Prompting
@@
3.2.1 Step 1: Entity Candidate Generation
Given a mention $m$ in document $D$, the first step is to find a subset of entity candidates $\mathcal{E}_{c}$ corresponding to $m$ within the knowledge base $\mathcal{E}$. We combine two strategies to generate entity candidates. First, we use statistical information of hyperlinks in the knowledge base to surface the most likely entity candidates $\mathcal{E}_{p}$ from $\mathcal{E}$ that exhibit syntactic similarity to the mentions within the document Ganea-EMNLP'17-deep_ed. However, this probabilistic approach, we call the Prior probability-approach or simply Prior, suffers from low recall because it only considers syntactic similarity. In low-recall cases, we additionally employ a dense retrieval model as a backup to augment the entity candidate generation. Specifically, we use the Bidirectional Encoder Representations from Transformers (BERT)-based entity linking (BLINK) model, which primarily uses cleaned Wikipedia hyperlinks to as its retrieval model Wu-EMNLP'20-BLINK, to generate extra entity candidates $\mathcal{E}_{r}$. Therefore, the final entity candidates set from step 1 is $\mathcal{E}_{c}=\mathcal{E}_{p}\cup\mathcal{E}_{r}$ from which we select the top-10 candidates.
3.2.2 Step 2: Augmentation by Prompting
@@
3.2.3 Step 3: Multiple-choice Selection by Prompting
@@
3.3 EntGPT-I: Entity Disambiguation with Instruction Tuning
In an initial study, described below, we found that EntGPT-P substantially reduces hallucination and helped to deliver more accurate results compared to non-augmented foundation models (see metrics in Tables 4 and  5). Nonetheless, EntGPT-P still under-performed supervised methods indicating room for improvement. To that end, we constructed an additional instruction-tuning model on top of generative LLMs through similar prompts as that of EntGPT-P. In this section, we discuss the processes involved in creating the EntGPT-I model including the construction of an instruction dataset and the instruction-tuning procedure.
3.3.1 Instruction Dataset Construction
An instruction dataset usually consists of three different elements: an instruction, an optional input, and an anticipated output ji2023survey. Usually in the form of natural language, an instruction provides the description of the task to the model, whereas any optional input can provide additional context. In the present work, we curate one of the largest manually annotated entity linking and entity disambiguation datasets, i.e., Accurate online Disambiguation of Entities in Text and Tables (AIDA) hoffart2011robust into the required format (instruction, input, output) needed to fine-tune the an LLM model. An example of such a training instance is shown in Fig. 1 (at right).
3.3.2 Instruction Tuning
We fine-tune two generative models (GPT3.5 and Llama2-chat) in a fully supervised manner on the curated instruction dataset. We used the OpenAI API platform kublik2023gpt to fine-tune GPT3.5, and we used Replicate API [^1] https://replicate.com/docs/get-started/python to fine-tune Llama2-chat. Fine-tuning was performed for only one epoch and we kept all other hyperparameters at their default settings. Although hyper-parameter tuning could potentially enhance performance, hyper-parameter search is outside the scope of this study. After fine-tuning, EntGPT-I is then evaluated on both the ED task and the question answering task.
4 Experiments
@@
  • 1.
    RQ1: Does multi-step prompting (EntGPT-P) and instruction tuning (EntGPT-I) produce more accurate entity disambiguation output?
  • 2.
@@
4.1 Datasets
Following general ED settings introduced in prior work Guo-Semantic_Web'16-Random_Walk; Kolitsas-CONLL'18-end2end; Ayoola-NAACL'22-ReFinED; Barba-ACL'22-Extend, we evaluated our models on six ED datsets: AIDA Milne-CIKM'08-Learning_to_Link, MSNBC (MSN) Cucerzan-EMNLP'07-MSNBC, Aquaint text corpus (AQUAINT/AQU) Milne-CIKM'08-Learning_to_Link, a multilingual training corpus ACE2004 (ACE04) Ratinov-ACL'11-Local_Global_Disambiguation, Walking Named Entity Disambiguation (WNED-WIKI) Guo-Semantic_Web'16-Random_Walk, and WNED-CWEB (CWEB) Gabrilovich-Dataset'13-Clueweb. We also consider five additional datasets from GERBIL Usbeck-WWW'15'-GERBIL: Keyphrase Overlap Relatedness for Entity Disambiguation (KORE50/KORE) Hoffart-CIKM'12-KORE, Open knowledge extraction challenge (OKE-2015/OKE15) Nuzzolese-SWEC'15-OKE15, OKE-2016 (OKE16) Nuzzolese-SWEC'16-OKE16, N3-Reuters-128 (REU) Roder-LREC'14-n3, and N3-RSS-500 (RSS) Roder-LREC'14-n3. We provide the statistics of these benchmarks, including the number of documents, and mentions in Table 1. Each of the experiments used Wikipedia as the background knowledge base. Following the example of Guo and Barbosa (Guo-Semantic_Web'16-Random_Walk), we manually removed spurious mentions that do not appear in the knowledge base as well as repeated documents and empty documents without any mentions.
We found that EntGPT-I (with GPT3.5 as the foundation model) outperformed other approaches on the average micro-F1 score (Tables 4 and  5), so we additionally evaluated EntGPT-I (GPT3.5) on six different commonsense reasoning (question answering) benchmarks: AI2 Reasoning Challenge-Challenge Set (ARC-C) clark2018think, OpenbbookQA (OBQA) mihaylov2018can, Abductive Natural Language Inference (aNLI) bhagavatula2019abductive, CommonsenseQA (CSQA) talmor2018commonsenseqa, Physical Interaction QA (PIQA) bisk2020piqa, and Social Interaction QA (SIQA) sap2019socialiqa.
4.2 Comparison Models